# Classification of $f(R)$ Theories Of Inflation

### v1

@@ 8 Conclusions
We have studied inflationary cosmology in the context of $f(R)$ theories. The vaious models can be classified by means of a mathematical analogy between the evolution of the background metric during inflation and the renormalization-group flow of quantum field theory. All the results are computed in the Jordan frame and expressed as expansions in a small parameter $\alpha$. A key quantity is the function $\beta_{\alpha}$, which is a power series in $\alpha$ that must be at least quadratic in order to realize slow-roll inflation. The classification is determined by the first nonvanishing term of $\beta_{\alpha}.$ This method helps us in identify the models that cannot explain the cosmic microwave background data. In particular, we have shown that only two classes of models can satisfy the experimental constraints. Among them, the Starobinsky model stands out as the only $f(R)$ polynomial extension of general relativity that can realize slow-roll inflation. All other polynomial $f(R)$ belong to a special class of models with a linear $\beta_{\alpha}$. The consequence of this property is that they can only realize constant-roll inflation. Since Starobinsky model and any of its polynomial extension belong to two different classes they cannot be analytically related in their parameters space not even perturbatively. The reason is that two competing expansions should be considered: one in the slow-roll parameter $\alpha$ and one in the coefficient $\bar{b}$ of the linear term in the beta function.
The classification presented here gives general NLL and NNLL formulas for the tensor and scalar power spectra, respectively, that can be used for any $f(R)$ model that belongs to class I or II, once the first few coefficient of $\beta_{\alpha}$ are derived. On a more speculative side, our results points out that principles typically used in particle physics could be relevant for selecting out theories that are suitable for primordial cosmology, given the fact that, once again, the Starobinsky model emerges for its uniqueness.

### v2

@@ 8 Conclusions
We have studied inflationary cosmology in the context of $f(R)$ theories. The various models can be classified by means of a mathematical analogy between the evolution of the background metric during inflation and the renormalization-group flow of quantum field theory. All the results are computed in the Jordan frame and expressed as expansions in a small parameter $\alpha$. A key quantity is the function $\beta_{\alpha}$, which is a power series in $\alpha$ that must be at least quadratic in order to realize slow-roll inflation. The classification is determined by the first nonvanishing term of $\beta_{\alpha}.$ This method helps us in identify the models that cannot explain the cosmic microwave background data. In particular, we have shown that only two classes of models can satisfy the experimental constraints. Among them, the Starobinsky model stands out as the only $f(R)$ polynomial extension of general relativity that can realize slow-roll inflation. All other polynomial $f(R)$ belong to a special class of models with a linear $\beta_{\alpha}$. The consequence of this property is that they can only realize constant-roll inflation. Since the Starobinsky model and any of its polynomial extension belong to two different classes they cannot be analytically related in their parameter space not even perturbatively. The reason is that two competing expansions should be considered: one in the slow-roll parameter $\alpha$ and one in the coefficient $\bar{b}$ of the linear term in the beta function.
The classification presented here gives general NLL and NNLL formulas for the tensor and scalar power spectra, respectively, that can be used for any $f(R)$ model that belongs to class I or II, once the first few coefficients of $\beta_{\alpha}$ are derived. On a more speculative side, our results point out that principles typically used in particle physics could be relevant for selecting out theories that are suitable for primordial cosmology, given the fact that, once again, the Starobinsky model emerges for its uniqueness.